Jump to main content or area navigation.

Contact Us

Waste Site Cleanup & Reuse in New England

Atlas Tack Corporation Superfund Site
Weekly Remedial Activities Update
Report # 14: October 9 - October 13, 2006

« Back to Atlas Tack Superfund Site Fact Sheet

1. ACTIVITIES PERFORMED DURING THE WEEK OF OCTOBER 9 - OCTOBER 13

  • Observed Columbus Day Holiday
  • Acquired equipment and expendable items (without costs) from the EPA/CENAE Norwood, MA project
  • Began the Transport and Disposal (T&D) of  RCRA Subtitle-D soil from stockpiles on Friday 13 October 2006
  • Attempted to complete the excavation in the East Fill.  Due to tidal influence, excavation activities were limited.
  • Construct an additional 240-linear foot berm along the northeast limits of the East Fill Area to prevent tidal waters from entering the excavation
  • Collect confirmatory soil samples from excavations in the East Fill Area
  • Collect waste characterization soil samples from soil originating from East Fill excavation for disposal profiling for off site landfill disposal
  • Continue to screen soil originating from East Fill Area
  • Continue to water the haul roads for dust control
  • Continue to collect  pre and post excavation survey data of the excavations
  • Global Positioning System (GIS) Surveying - Trimble-RTK GIS representative upgrades Trimble GIS software on 09 October 2006
  • Continue water handling activities
  • Continue with air monitoring program
  • Completed  2nd round of monitoring well water sampling

2. PROJECTED ACTIVITIES FOR THE WEEK OF OCTOBER 16 - OCTOBER 20

  • Continue to T&D RCRA Subtitle D soil from stockpiles (2,500-cy)
  • Complete landfill approval process of  additional Subtitle-D soil
  • Perform re-excavation activities in the East Fill Area (if applicable)
  • re-excavation of soils along the western boundary of the CID
  • Collect confirmatory soil samples from excavations in the East Fill Area
  • Collect waste characterization soil samples from soils originating from East Fill excavation for disposal profiling for off site landfill disposal
  • Continue to screen soils originating from East Fill Area
  • Continue to water the haul roads for dust control
  • Continue to collect  pre and post excavation survey data of the excavations
  • Continue water handling activities
  • Continue with air monitoring program

3. AIR MONITORING INFORMATION

WESTON monitored for particulate (dust) during intrusive activities performed during this reporting period. WESTON recognizes EPA's level of concern for "PM10" to caution sensitive groups, including children, the elderly, and people with heart or lung disease, about exposure to coarse dust particles. The PM10 level of concern for sensitive groups is 150 µg/m3 for 24-hour average exposure.

Perimeter Air Monitoring is performed at up to (6) locations throughout a given work day. Readings are collected at 60 second intervals throughout the day and reported as a daily average.

If exceedances are identified, an increased effort of engineering controls (water suppression) will be utilized to adhere to the PM10 level of concern. If engineering controls appear to be inadequate, any dust generating activities will stop. Current dust control practices will be re-evaluated until concentrations are reduced to below the PM10 standard.

Results for 5 October – 11 October 2006 can be reviewed below.

Weekly Dust Monitoring Summary
Week #14: Results for Thursday through Wednesday**

The EPA's PM10 level of concern for sensitive groups is 150 µg/m3 for 24-hour average exposure

Date Pleasant Street (µg/m3) Adams Street (µg/m3) Tripp Street (µg/m3) Egypt Road (µg/m3) Hurricane Barrier (µg/m3) Church Street 2 (µg/m3)
9/28/2006 1 3 6 56 0 0
9/29/2006 0 0 7 NSC 0 0
10/2/2006 Holiday Holiday Holiday Holiday Holiday Holiday
10/3/2006 0 0 0 NSC 22 0
10/4/2006 6 6 4 0 0 8

Notes:
All readings reflect the daily average calculated over the course of the workday.
All readings collected with a Thermo Electron PDR 1000.
NSC: No sampling conducted; No intrusive work conducted in that area.
Precip: No sampling conducted due to precipitation.
*: Monitoring conducted over 24-hour period.
**: No work and air monitoring occurs on weekend days

4. QUANTITY SUMMARY

Report # 14 - SUMMARY
TO DATE
ACTIVITIES TOTAL QUANTITY TOTAL QUANTITY
C.I.D. Excavation (Re-Digs) None 6,304 CY
East Fill Excavation 500 CY 8,224 CY
Soils T&D to Crapo Landfill None 1,237 Tons
Soils T&D to Marion Landfill None 1,940 Tons

Notes:
CY = volume of soil/debris in cubic yards

Jump to main content.